The prelim2e package^{∗}

Martin Schröder

Friedrich-Humbert-Straße 124 D-28759 Bremen MS@Dream.HB.North.DE (INTERNET)

1995/01/26

Abstract

This package allows the marking of (preliminary) versions of a document on the output.

Contents

1 Introduction 1

2 Usage 1 3 Options 1

4 Required packages 2

1 Introduction

This package allows the marking of (preliminary) versions of a document. This is achieved using the command \PrelimText, whose expansion is added below the footer of every page of a document (look at the bottom of this page for an example).

2 Usage

Simply using this package via \usepackage{prelim} does produce a text in the form of “Preliminary version – January 4, 1996 – 16:19”.

\PrelimText The text is produced by the command \PrelimText, which can be changed via \renewcommand* or by setting options at the \usepackage command (see section 3).

3 Options

The package has the following options:

draft If this option is used a text appears below the normal pagebody. It is the default.

final If this option is used prelim2e does produce no text.

1 english This sets the text to “Preliminary version”. It is the default. german This sets the text to “Vorläufige Version”. It does not use the german or

∗The version number of this file is v1.00, last revised 1995/01/26.

prelim2e package – Version v1.00 – Documentation L^{A}TEXed on January 4, 1996 at 16:19

babel package.

4 Required packages

The package requires the following packages: everyshi This package is used to implement the setting of the text below the normal

pagebody. printtim This package is used to typeset the current time.

prelim2e package – Version v1.00 – Documentation L^{A}TEXed on January 4, 1996 at 16:19